OpenStudy (anonymous):

Forty-two is what percent of 70? A. 21% B. 60% C. 74% D. 167%

OpenStudy (anonymous):

42/70 = 6/10 = 60/100 = 60% ;)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Which word is used to name the bold numbers and variables? 2x + 5 = 17 A. minuend B. addend C. dividend D. divisor